# An Information-Theoretic Approach to Understanding Transformers’ In-Context Learning of Variable-Order Markov Chains

### 28 Mar 2026 v2

@@ G.2.1 Finite-memory context-extension layer
Take $c=\infty$ or sufficiently large. It is seen that the $m$-th head essentially copies the $m$-th earlier symbol to stack at the $(m+1)$-th position below the original symbol $\vec{x}_{i}$. Together with the residual link, the attention layer gives exactly the $\vec{h}^{(2)}_{i}$ shown in (29) while the FF layer in this layer can be set to all zeros.
@@ G.2.2 Statistics collection layer
Since $p_{i,s_{i,l}}^{e}$ and $\vec{p}_{i,s_{i,l}}$ in (13) are simple functions of $\vec{n}_{i,s_{i,l}}$, they can be approximated by a sufficiently wide FF layer, yielding the following output
@@ G.2.3 Inductive CTW layer
There exists a $A$-head transformer layer that can perform the induction: Takes $\vec{H}^{(\ell)}$ in (42) as input and outputs $\vec{H}^{(\ell+1)}$. And the final readout layer taking $\vec{H}^{(D+2)}$ as input can output the $A$-dimensional Bayesian optimal next token prediction vector $P_{\pi_{\text{CTW}}}(\cdot|x_{1-D}^{n})=\sum_{l=0,\ldots,D}~\omega_{n,l}\vec{p}_{n,s_{n,l}}.$
@@
For any fixed $\ell=3,4,\ldots,2+D$, we specify the construction for the $\ell$-th transformer layer. It contains $A$ heads and for each $m=1,2,\ldots,A$, the $Q,K,V$ matrices are
@@
where $\vec{e}_{m}$ is the $A$-dimensional one-hot vector at position $m$, and $\text{place}_{\ell}=(M^{(1)}+D+2)A+D+\ell-1$ is the index of element $\ell^{w}_{i,s_{i,D+3-\ell}}$ in $\vec{h}_{i}^{(\ell)}$. The corresponding query, key, and value vectors after the mapping are
@@
Recall $\ln(\omega_{n,l})-\ln(\omega_{n,l-1})=\ln(1-\lambda)-\mathbb{I}_{l=D}\ln(\lambda)+\ell^{e}_{n,s_{n,l}}-\ell^{e}_{n,s_{n,l-1}}+\sum_{q\in\mathcal{A}}\ell^{w}_{n,qs_{n,l-1}}-\ell^{w}_{n,s_{n,l}}$ by Theorem 2. $\delta_{i,D+3-\ell}=\ln(\omega_{i,D+3-\ell})-\ln(\omega_{i,D+2-\ell})$ can be computed by $\vec{a}_{i}^{(\ell)}$ and thus $\vec{h}_{i}^{(\ell+1)}$ can be conveniently computed via the FF layer following the $\ell$-th multi-head attention layer.
The final readout layer approximates an $A$-dimensional vector
@@
by an FF layer taking input

### 31 Mar 2026 v3

@@ G.2.1 Finite-memory context-extension layer
Take $c=\infty$ or sufficiently large. It is seen that the $m$-th head essentially copies the $m$-th earlier symbol to stack at the $(m+1)$-th position below the original symbol $\vec{x}_{i}$. Together with the residual link, the attention layer gives exactly the $\vec{h}^{(2)}_{i}$ shown in (29) while the $F$ layer in this layer can be set to all zeros.
@@ G.2.2 Statistics collection layer
Since $p_{i,s_{i,l}}^{e}$ and $\vec{p}_{i,s_{i,l}}$ in (13) are simple functions of $\vec{n}_{i,s_{i,l}}$, they can be approximated by a sufficiently wide $F$ layer, yielding the following output
@@ G.2.3 Inductive CTW layer
There exists a AA-head transformer layer that can perform the induction: Takes $\vec{H}^{(\ell)}$ in (42) as input and outputs $\vec{H}^{(\ell+1)}$. And the final readout layer taking $\vec{H}^{(D+2)}$ as input can output the AA-dimensional Bayesian optimal next token prediction vector $P_{\pi_{\text{CTW}}}(\cdot|x_{1-D}^{n})=\sum_{l=0,\ldots,D}~\omega_{n,l}\vec{p}_{n,s_{n,l}}.$
@@
For any fixed $\ell=3,4,\ldots,2+D$, we specify the construction for the $\ell$-th transformer layer. It contains AA heads and for each $m=1,2,\ldots,A$, the $Q,K,V$ matrices are
@@
where $\vec{e}_{m}$ is the AA-dimensional one-hot vector at position $m$, and $\text{place}_{\ell}=(M^{(1)}+D+2)A+D+\ell-1$ is the index of element $\ell^{w}_{i,s_{i,D+3-\ell}}$ in $\vec{h}_{i}^{(\ell)}$. The corresponding query, key, and value vectors after the mapping are
@@
Recall $\ln(\omega_{n,l})-\ln(\omega_{n,l-1})=\ln(1-\lambda)-\mathbb{I}_{l=D}\ln(\lambda)+\ell^{e}_{n,s_{n,l}}-\ell^{e}_{n,s_{n,l-1}}+\sum_{q\in\mathcal{A}}\ell^{w}_{n,qs_{n,l-1}}-\ell^{w}_{n,s_{n,l}}$ by Theorem 2. $\delta_{i,D+3-\ell}=\ln(\omega_{i,D+3-\ell})-\ln(\omega_{i,D+2-\ell})$ can be computed by $\vec{a}_{i}^{(\ell)}$ and thus $\vec{h}_{i}^{(\ell+1)}$ can be conveniently computed via the $F$ layer following the $\ell$-th multi-head attention layer.
The final readout layer approximates an AA-dimensional vector
@@
by an $F$ layer taking input